If the measure of angle RQT=155 degrees what is the measure of angle TQS
R________Q\______/T that is the best way I can draw it the forward slash by Q shows the way that ray QS is going the back slash is the way that QT is going in between RQS is the equation 4x-20 and in between TQS is the equation 3x+14 i've been trying to help my daughter but i can't figure it out

Answer by stanbon(75887)

If the measure of angle RQT=155 degrees what is the measure of angle TQS
R________Q\______/T that is the best way I can draw it the forward slash by Q shows the way that ray QS is going the back slash is the way that QT is going in between RQS is the equation 4x-20 and in between TQS is the equation 3x+14 i've been trying to help my daughter but i can't figure it out
------
angleRQT = angleRQS + angleTQS
155 = 4x-20 + 3x+14
------
7x - 6 = 155
7x = 161
x = 23
-----
angleRQS = 4x-20 = 4*23-20 = 92-20 = 72 degrees
angleTQS = 3x+14 = 3*23+14 = 69+14 = 83 degrees
--------------
Cheers,
Stan H.